# Asymptotically Optimal Tests for One- and Two-Sample Problems

### v1

@@ I Introduction
The one-sample test, often referred to as the universal hypothesis testing in the statistical literature, was formulated and solved in Hoeffding’s seminal work [7]. Hoeffding showed that the so-called likelihoood ratio test achieves the optimal type II error exponent, which is the relative entropy between the nominal distribution and the unknown true distribution under $H_{1}$. Hoeffding’s test is a threshold test of the likelihood ratio between the empirical distribution and the nominal distribution. The test statistic reduces to the simple form of the relative entropy between the empirical distribution and the nominal distribution. Through the use of a properly controlled vanishing threshold, Hoeffding showed that the test is asymptotically optimal in that it achieves the optimal type II error exponent subject to a level constraint on the type I error. Hoeffding’s work inspired a large body of follow-up research, including attempts to extend the solution to continuous distributions [13, 12].
The two-sample test, also called homogeneity testing [8] or out-of-distribution sample detection [6], has found applications in various fields [2, 9]. While the two-sample test has been studied in the statistical literature, it has attracted significant attention recently in the machine learning community thanks to the discovery of the maximum mean discrepancy (MMD) test [5] that has shown promising performance across various datasets. This has inspired extensive follow-up research; and it was recently established that MMD with suitable kernels was indeed asymptotically optimal in that it achieves the optimal type II error exponent [15, 14] when subject to a level constraint on the type I error.
In this work, we revisit both the one-sample and two-sample problems for distributions with finite support. A more streamlined achievability proof of Hoeffding’s likelihood ratio test is provided that also offers a clean interpretation. The proof, as well as the test statistic itself, naturally extends to the two-sample test. More precisely, the threshold test of the relative entropy between the two empirical distributions is shown to achieve the optimal error exponent for the two-sample test, which coincides with Renyi-divergence with $\alpha=0.5$ when the two sequences are of identical length. Finally, a strong converse for the two-sample test is obtained, i.e., any test whose type II error probability decays faster than the optimal type II error exponent will have its type I error probability approaching $1$.
@@ II-A Notation and Preliminaries
Take $d]=\{1,2,\dots,d\}$, $\log$ denotes base 2 log, $\ln$ denotes log base $e$, and we use $\operatorname{supp}(P)$ to describe the support of a probability distribution $P$. For two probability distributions $P,Q$ , $P<<Q$ is short hand for $\operatorname{supp}(P)\subseteq\operatorname{supp}(Q)$. We use $X^{n}=(X_{1},\dots,X_{n})$ to refer to a collection of $n$ samples. We say $X^{n}\sim Q$ to denote $(X_{1},\dots,X_{n})$ being independent and identically distributed samples taken from the distribution $Q$. Let $\mathcal{P}^{d}$ be the probability simplex over $d$ elements such that
@@
For $P\in\mathcal{P}^{d}$, we use $p_{\min}$ to denote the minimum non-zero probability in $P$. Denote the total variation distance, Kullback-Leibler divergence (KLD) between distributions $P,Q\in\mathcal{P}^{d}$ respectively by
$\operatorname{TV}(P_{0},P_{1})=\frac{1}{2}\left\|P_{0}-P_{1}\right\|_{1}$ (3)
$D\left(P\|Q\right)=\sum_{i=1}^{d}p_{i}\log\frac{p_{i}}{q_{i}},$ (4)
@@
$\mathcal{B}(P,c)$ $=\{F\in\mathcal{P}^{d}:D\left(F\|P\right)\leq c\},$ (6)
and we use $\mathcal{B}^{C}(P,c)$ to denote its complement in $\mathcal{P}^{d}$. For $X^{n}\sim P$ and $Y^{m}\sim Q$, denote the empirical distributions of these samples as $\hat{P}_{X^{n}}$ and $\hat{Q}_{Y^{n}}$. Throughout this paper, we will make heavy use of the following large deviation result [4, Theorem 12.4.1]
@@
$\mathbb{P}[\hat{P}_{X^{n}}\in E]\leq(n+1)^{d}2^{-nD\left(F^{*}\|P\right)},$ (7)
where $P^{*}=\operatorname*{arg\,min}_{F\in E}D\left(F\|P\right)$.
@@
$\left|D\left(A\|C\right)-D\left(B\|C\right)\right|$ $\leq 2\log\left(\frac{1}{c_{\min}}\right)\varepsilon+\frac{2\sqrt{2}}{\ln(2)}\sqrt{\varepsilon}$ (8)
$\triangleq g(C,\varepsilon).$ (9)
@@
$\lim_{n\to\infty}\frac{1}{n}\log\beta^{*}_{n}(\varepsilon)=D\left(P_{0}\|P_{1}\right).$ (11)
@@ II-B One-Sample Problem
Figure 1: One-Sample Problem
@@
where $c_{n}n^{2}\to\infty$. Hoeffding refers to this as the likelihood ratio test as the test statistic, as $D(\hat{Q}_{X^{n}}\|P)$ is equivalent to the log likelihood ratio between the empirical distribution and the nominal distribution. He showed that if $c_{n}\to 0$, this likelhood ratio test, herein denoted by $T_{\text{Hoff},n}$, has the following probabilities:
@@
$\beta(T_{\text{Hoff},n})$ $=2^{-nD\left(P\|Q\right)+O(\log n)},$ (14)
@@
$\liminf_{n\to\infty}\frac{1}{n}\log\beta(T_{\text{Hoff},n})\geq D\left(P\|Q\right).$ (15)
Combined with Theorem˜2, this establishes the fact that the test $T_{\text{Hoff},n}$ is asymptotically optimal with type II error exponent $D\left(P\|Q\right)$.
@@ II-C Two-Sample Problem
The two-sample test involves samples $X^{n}\sim P$, and $Y^{n}\sim Q$, with both $P$ and $Q$ unknown. The task is to determine if both samples are generated from the same distribution, and the hypothesis test is again in exactly the same form as (1). Figure˜2 gives a visual description of this problem.
Figure 2: Two-Sample Problem
@@
$\lim_{n\to\infty}\frac{1}{n}\log\beta^{*}_{n}(\varepsilon)=\inf_{F\in\mathcal{P}^{d}}D\left(F\|P\right)+D\left(F\|Q\right).$ (16)
@@ III-A One Sample Test
$D\left(F\|Q\right)\approx D\left(P\|Q\right),\quad\forall F\in\mathcal{B}(P,c_{n}).$ (17)
@@
Further, if $D\left(P\|Q\right)=+\infty$, then there exists $N>0$ such that for any $n\geq N$,
@@
If $D\left(P\|Q\right)<+\infty$, then instead we get
$\lim_{n\to\infty}\frac{1}{n}\log\beta(A_{n})=D\left(P\|Q\right).$ (20)
The proof can be found in Section˜IV-A.
While the optimality property was originally established in [7], the proof, which offers a simple geometric interpretation, allows one to directly contruct a test for the two-sample problem with the desired optimality.
@@ III-B Two Sample Test
Recall that $\hat{P}_{X^{n}}$ and $\hat{Q}_{Y^{n}}$ denote respectively the empirical distributions of $X^{n}$ and $Y^{n}$. The one-sample test uses the KLD ball around the nominal distribution with a vanishing radius as the acceptance region. The rate with which it vanishes is controlled in a manner to ensure type I error constraint is satisfied, i.e., the probability that the empirical distributions fall within the KLD ball should be sufficiently large. The same geometric interpretation can be used in constructing a test that works for the two-sample setting. As the sample size increases, the empirical distributions will concentrate around the true distribution, i.e., $\hat{P}_{X^{n}}$ will likely fall within the KLD ball centered at $P$ whereas $\hat{Q}_{Y^{n}}$ will likely fall within the KLD ball centered at $Q$. If $P=Q$, then the KLD ball will overlap and one can resort to Bonferroni inequality to argue that $\hat{P}_{X^{n}}$ will concentrate around the KLD ball centered at $\hat{Q}_{Y^{n}}$ and vice versa. This motivates the following test for the two-sample problem.
$\hat{H}=\begin{cases}H_{0}&:D\left(\hat{P}_{X^{n}}\|\hat{Q}_{Y^{n}}\right)\leq c_{n},\\ H_{1}&:D\left(\hat{P}_{X^{n}}\|\hat{Q}_{Y^{n}}\right)>c_{n},\end{cases}$ (21)
where $c_{n}=\Theta(n^{-1/2})$ is chosen to satisfy a type I error constraint***Our choice of $c_{n}=\Theta(n^{-1/2})$ is arbitrary, the result holds for $c_{n}=n^{-1+\gamma}$ for any $\gamma\in(0,1)$.. The following theorem states that this test achieves the optimal type II error exponent.
@@
Using the test described in Equation˜21 as $A_{n}$, we have, $\forall P\in\mathcal{P}_{d}$,
$\limsup_{n\to\infty}\mathbb{P}\left[(\hat{P}_{X^{n}},\hat{P}_{Y^{n}})\in A_{n}^{C}|X^{n},Y^{n}\sim P\right]\leq\varepsilon.$ (22)
@@
$\mathbb{P}\left[(\hat{P}_{X^{n}},\hat{Q}_{Y^{n}})\in A_{n}|(X^{n},Y^{n})\sim(P,Q)\right]=0.$ (23)
@@
$\begin{split}-\liminf_{n\to\infty}\frac{1}{n}&\log\mathbb{P}\left[(\hat{P}_{X^{n}},\hat{Q}_{Y^{n}})\in A_{n}|(X^{n},Y^{n})\sim(P,Q)\right]\\ &\geq\left[D\left(F^{*}\|P\right)+D\left(F^{*}\|Q\right)\right].\end{split}$ (24)
where $F^{*}=\operatorname*{arg\,min}_{F\in\mathcal{P}^{d}}D\left(F\|P\right)+D\left(F\|Q\right)$.
A proof sketch can be found in Section˜IV-B.
Notice that the test described in Equation˜21 is not symmetric in $P$ and $Q$, while the problem itself is, both in the problem formulation and in the optimal error exponent. Indeed, the result holds when replacing $D\left(\hat{P}_{X^{n}}\|\hat{Q}_{Y^{n}}\right)$ with $D\left(\hat{Q}_{Y^{n}}\|\hat{P}_{X^{n}}\right)$, and or the minimum of the two. Using the minimum test statistic yields a better type I error while achieving the same asymptotic type II exponent.
@@
For a given test $T_{n}$, if for a particular $P,Q\in\mathcal{P}^{d}$ with $\operatorname{supp}(P)\cap\operatorname{supp}(Q)\neq\emptyset$ and $P\neq Q$, we have
$\begin{split}-&\liminf_{n\to\infty}\frac{1}{n}\log\mathbb{P}\left[(\hat{P}_{X^{n}},\hat{Q}_{Y^{n}})\in T_{n}|(X^{n},Y^{n})\sim(P,Q)\right]\\ &>\left[D\left(F^{*}\|P\right)+D\left(F^{*}\|Q\right)\right].\end{split}$ (25)
@@
$\lim_{n\to\infty}\mathbb{P}\left[(\hat{P}_{X^{n}},\hat{Q}_{Y^{n}})\not\in T_{n}|X^{n},Y^{n}\sim F^{*}\right]=1.$ (26)
A sketch of the proof can be found in Section˜IV-C.

### v2

@@ I Introduction
The one-sample test, often referred to as the universal hypothesis testing in the statistical literature, was formulated and solved in Hoeffding’s seminal work [9]. Hoeffding showed that the so-called likelihood ratio test achieves the optimal type II error exponent, which is the relative entropy between the nominal distribution and the unknown true distribution under $H_{1}$. Hoeffding’s test is a threshold test of the likelihood ratio between the empirical distribution and the nominal distribution. The test statistic reduces to the simple form of the relative entropy between the empirical distribution and the nominal distribution. Through the use of a properly controlled vanishing threshold, Hoeffding showed that the test is asymptotically optimal in that it achieves the optimal type II error exponent subject to a level constraint on the type I error. Hoeffding’s work inspired a large body of follow-up research, including attempts to extend the solution to continuous distributions [18, 17]. A second order analysis, which describes finite sample behavior of the error, was carried out for one-sample tests [6].
The two-sample test, also called homogeneity testing [10] or out-of-distribution sample detection [8], has found applications in various fields [2, 12]. While the two-sample test has been studied in the statistical literature, it has attracted significant attention recently in the machine learning community thanks to the discovery of the maximum mean discrepancy (MMD) test [4] that has shown promising performance across various datasets. This has inspired extensive follow-up research; and it was recently established that MMD with suitable kernels was indeed asymptotically optimal in that it achieves the optimal type II error exponent [21, 20] when subject to a level constraint on the type I error. A closely related formulation is the Gutman classification problem[5], where training sequences $X^{n}\sim P_{1}$ and $Y^{n}\sim P_{2}$ are given, along with a test sequence $Z^{m}\sim Q$. The goal is to determine if $Q=P_{1}$ or $Q=P_{2}$, using only the empirical samples. In [19] a second order asymptotic analysis was performed for the Gutman classication problem. At the same time as this work was released, a second order analysis of two-sample hypothesis testing was executed in [7].
In this work, we revisit both the one-sample and two-sample problems for distributions with finite support. A more streamlined achievability proof of Hoeffding’s likelihood ratio test is provided that also offers a clean interpretation. The proof, as well as the test statistic itself, naturally extends to the two-sample test. More precisely, the threshold test of the relative entropy between the two empirical distributions is shown to achieve the optimal error exponent for the two-sample test, which coincides with Renyi-divergence with $\alpha=\frac{1}{2}$ when the two sequences are of identical length. Finally, a strong converse for the two-sample test is obtained, i.e., any test whose type II error probability decays faster than the optimal type II error exponent will have its type I error probability approaching $1$.
@@ II-A Notation and Preliminaries
Take $d]=\{1,2,\dots,d\}$, $\log$ denotes base 2 logarithm, $\ln$ denotes logarithm base $e$, and we use $\operatorname{supp}(P)$ to describe the support of a probability distribution $P$. For two probability distributions $P,Q$ , $P<<Q$ is short hand for $\operatorname{supp}(P)\subseteq\operatorname{supp}(Q)$. We use $X^{n}=(X_{1},\dots,X_{n})$ to refer to a collection of $n$ samples. We say $X^{n}\sim Q$ to denote $(X_{1},\dots,X_{n})$ being independent and identically distributed samples taken from the distribution $Q$. Let $\mathcal{P}^{d}$ be the probability simplex over $d$ elements such that
@@
For $P\in\mathcal{P}^{d}$, we use $p_{\min}$ to denote the minimum non-zero probability in $P$. Denote the total variation distance between distributions $P,Q\in\mathcal{P}^{d}$ by
$\operatorname{TV}(P,Q)=\frac{1}{2}\left\|P-Q\right\|_{1},$ (3)
and the Kullback-Leibler divergence (KLD)
$D\left(P\middle\|Q\right)=\sum_{i=1}^{d}p_{i}\log\frac{p_{i}}{q_{i}},$ (4)
@@
$\mathcal{B}(P,c)$ $=\{F\in\mathcal{P}^{d}:D\left(F\middle\|P\right)\leq c\},$ (6)
and we use $\mathcal{B}^{C}(P,c)$ to denote its complement in $\mathcal{P}^{d}$. For $X^{n}\sim P$ and $Y^{m}\sim Q$, denote the empirical distributions of these samples as $\hat{P}_{X^{n}}$ and $\hat{Q}_{Y^{n}}$. Throughout this paper, we will make heavy use of the following large deviation result [3, Theorem 12.4.1]
@@
$\mathbb{P}[\hat{P}_{X^{n}}\in E]\leq(n+1)^{d}2^{-nD\left(F^{*}\middle\|P\right)},$ (7)
where $F^{*}=\operatorname*{arg\,min}_{F\in E}D\left(F\middle\|P\right)$.
@@
$\left|D\left(A\middle\|C\right)-D\left(B\middle\|C\right)\right|$ $\leq 2\log\left(\frac{1}{c_{\min}}\right)\varepsilon+\frac{2\sqrt{2}}{\ln(2)}\sqrt{\varepsilon}$ (8)
$\triangleq g(C,\varepsilon),$ (9)
where $c_{\min}$ is the minimum non-zero probability in $C$.
@@
$\lim_{n\to\infty}-\frac{1}{n}\log\beta^{*}_{n}(\varepsilon)=D\left(P_{0}\middle\|P_{1}\right).$ (11)
@@ II-B One-Sample Problem
Figure 1: In the one-sample problem we know the nominal distribution $P$, and all information about $Q$ is taken from the sample $X^{n}$
@@
Hoeffding refers to this as the likelihood ratio test, as the test statistic, $D(\hat{Q}_{X^{n}}\|P)$, is equivalent to the log likelihood ratio between the empirical distribution and the nominal distribution. He showed that if $c_{n}\to 0$, this likelihood ratio test, herein denoted by $T_{\text{Hoff},n}$, has the following probabilities:
@@
$\beta(T_{\text{Hoff},n})$ $=2^{-nD\left(P\middle\|Q\right)+O(\log n)},$ (14)
@@
$\liminf_{n\to\infty}-\frac{1}{n}\log\beta(T_{\text{Hoff},n})\geq D\left(P\middle\|Q\right).$ (15)
Combined with Theorem˜2, this establishes the fact that the test $T_{\text{Hoff},n}$ is asymptotically optimal with type II error exponent $D\left(P\middle\|Q\right)$.
@@ II-C Two-Sample Problem
In the two-sample problem we are given samples $X^{n}\sim P$ and $Y^{m}\sim Q$, with both $P$ and $Q$ unknown. The task is to determine if both samples are generated from the same distribution, and the hypothesis test is in exactly the same form as (1). Here we allow $n\neq m$, but under we assume that $\lim_{n,m\to\infty}\frac{n}{n+m}=w\in\left[\frac{1}{2},1\right]$, representing potentially different sample rates for the two distributions. We make the assumption that $n>m$ without loss of generality, as we can simply exchange $X^{n}$ and $Y^{m}$ in our definition. Figure˜2 gives a visual description of this problem.
Figure 2: In the two-sample problem we have no knowledge of either true distribution, with all information about $P$ and $Q$ coming from the two samples $X^{n}$ and $Y^{m}$
For a given $n,m$, the type II error rate is
$\beta_{n,m}^{*}(\varepsilon)=\inf_{A_{n,m}}\{\beta(A_{n,m}):\alpha(A_{n,m})\leq\varepsilon\}$ (16)
@@
$\begin{split}\lim_{\begin{subarray}{c}n,m\to\infty\end{subarray}}\frac{-1}{n+m}&\log\beta_{n,m}^{*}(\varepsilon)\\ =\inf_{F\in\mathcal{P}^{d}}&wD\left(F\middle\|P\right)+(1-w)D\left(F\middle\|Q\right).\end{split}$ (17)
@@ III-A One Sample Test
$D\left(F\middle\|Q\right)\approx D\left(P\middle\|Q\right),\quad\forall F\in\mathcal{B}(P,c_{n}).$ (18)
@@
Further, if $D\left(P\middle\|Q\right)=+\infty$, then there exists $N>0$ such that for any $n\geq N$,
@@
If $D\left(P\middle\|Q\right)<+\infty$, then instead we get
$\lim_{n\to\infty}-\frac{1}{n}\log\beta(A_{n})=D\left(P\middle\|Q\right).$ (21)
The proof can be found in Appendix˜B .
While the optimality property was originally established in [9], our proof, which offers a simple geometric interpretation, allows one to directly construct a test for the two-sample problem with the desired optimality.
@@ III-B Two Sample Test
Recall that $\hat{P}_{X^{n}}$ and $\hat{Q}_{Y^{m}}$ denote respectively the empirical distributions of $X^{n}$ and $Y^{m}$. The one-sample test uses the KLD ball around the nominal distribution with a vanishing radius as the acceptance region. The rate with which it vanishes is controlled in a manner to ensure type I error constraint is satisfied, i.e., the probability that the empirical distributions fall within the KLD ball should be sufficiently large. The same geometric interpretation can be used in constructing a test that works for the two-sample setting. As the sample size increases, the empirical distributions will concentrate around the true distribution, i.e., $\hat{P}_{X^{n}}$ will likely fall within the KLD ball centered at $P$ whereas $\hat{Q}_{Y^{m}}$ will likely fall within the KLD ball centered at $Q$. If $P=Q$, then the KLD ball will overlap and one can resort to Bonferroni inequality to argue that $\hat{P}_{X^{n}}$ will concentrate around the KLD ball centered at $\hat{Q}_{Y^{m}}$ and vice versa. This motivates the following test for the two-sample problem.
$\hat{H}=\begin{cases}H_{0}&:D\left(\hat{P}_{X^{n}}\middle\|\hat{Q}_{Y^{m}}\right)\leq c_{m},\\ H_{1}&:D\left(\hat{P}_{X^{n}}\middle\|\hat{Q}_{Y^{m}}\right)>c_{m},\end{cases}$ (22)
where $c_{m}=\Theta(m^{-1/2})$ is chosen to satisfy a type I error constraint***Our choice of $c_{m}=\Theta(m^{-1/2})$ is arbitrary, the result holds for $c_{m}=m^{-1+\gamma}$ for any $\gamma\in(0,1)$.. The following theorem states that this test achieves the optimal type II error exponent.
@@
Given $\varepsilon\in(0,1)$ the constraint on the type I error, using the test described in Equation˜22 as $A_{n,m}$, we have, $\forall P\in\mathcal{P}_{d}$,
$\limsup_{n\to\infty}\mathbb{P}\left[(\hat{P}_{X^{n}},\hat{P}_{Y^{m}})\in A_{n,m}^{C}|X^{n},Y^{m}\sim P\right]\leq\varepsilon.$ (23)
@@
$\mathbb{P}\left[(\hat{P}_{X^{n}},\hat{Q}_{Y^{m}})\in A_{n,m}|(X^{n},Y^{m})\sim(P,Q)\right]=0.$ (24)
@@
$\begin{split}\liminf_{n,m\to\infty}-\frac{1}{n+m}&\log\mathbb{P}\left[(\hat{P}_{X^{n}},\hat{Q}_{Y^{m}})\in A_{n.m}|(X^{n},Y^{m})\sim(P,Q)\right]\\ &\geq\left[wD\left(F^{*}\middle\|P\right)+(1-w)D\left(F^{*}\middle\|Q\right)\right].\end{split}$ (25)
where $F^{*}=\operatorname*{arg\,min}_{F\in\mathcal{P}^{d}}wD\left(F\middle\|P\right)+(1-w)D\left(F\middle\|Q\right)$.
A proof sketch can be found in Appendix˜C .
Notice that the test described in Equation˜22 is not symmetric in $P$ and $Q$, while the problem itself is, both in the problem formulation and in the optimal error exponent. Indeed, the result holds when replacing $D\left(\hat{P}_{X^{n}}\middle\|\hat{Q}_{Y^{m}}\right)$ with $D\left(\hat{Q}_{Y^{m}}\middle\|\hat{P}_{X^{n}}\right)$, and or the minimum of the two. Using the minimum test statistic yields a better type I error while achieving the same asymptotic type II exponent.
@@
For a given test $T_{n,m}$, if for a particular $P,Q\in\mathcal{P}^{d}$ with $\operatorname{supp}(P)\cap\operatorname{supp}(Q)\neq\emptyset$ and $P\neq Q$, we have
$\begin{split}&\liminf_{n,m\to\infty}-\frac{1}{n+m}\log\mathbb{P}\left[(\hat{P}_{X^{n}},\hat{Q}_{Y^{m}})\in T_{n,m}|(X^{n},Y^{m})\sim(P,Q)\right]\\ &>\left[wD\left(F^{*}\middle\|P\right)+(1-w)D\left(F^{*}\middle\|Q\right)\right].\end{split}$ (26)
@@
$\lim_{n\to\infty}\mathbb{P}\left[(\hat{P}_{X^{n}},\hat{Q}_{Y^{m}})\not\in T_{n,m}|X^{n},Y^{m}\sim F^{*}\right]=1.$ (27)
The proof can be found in Appendix˜D